ABC Ltd issues a $20 million IPO providing proceeds to ABC of $4.5 per share, from an offer price to the public of $5 per share. The company's legal fees, ASIC registration fees, and other administrative costs are $314,000. The company's share price increases 8 per cent on the first day. What is the underpricing cost to the company of issuing the securities? (in millions of dollars to the nearest three decimal places; don’t use the $ sign eg 7.897)?
Answer #1

Under pricing cost is the difference between the market price of shares on first day end and the offer price.

No. of shares issued = gross amount issued / offer price = 20,000,000 / 5 = 4,000,000 shares issued.

Under pricing cost = No of shares issued * ( first day ending price - offer price )

First day ending price = 5 + 5 *8% = $ 5.4 per share

Underpricing cost = 4,000,000 * ( 5.4 - 5)

Underpricing cost = $ 1,600,000

Answer Format is 1.600